ah42632:

ABCD is an isosceles trapezoid with legs Modifying above upper A upper B with barand modifying above upper C upper D with barand base modifying above B C with bar. If the length ofModifying above upper A upper B with baris 7y – 4, the length of modifying above B C with bar is 4y – 6 and the length of modifying above upper C upper D with baris 8y – 18, find the value of y.

CHEERQUEEN:

C i am prettie sure

ah42632:

its only an open writing type of question, it doesn´t show answers:(

CHEERQUEEN:

oh ummmmmm hang on

CHEERQUEEN:

do u have any gusee

CHEERQUEEN:

We are told that the legs of the isosceles trapezoid are AB and CD AB = 7y - 4 CD = 8y - 18 In any isosceles trapezoid, the two legs are of equal length. Therefore, AB = CD ⇒ 7y - 4 = 8y - 18 Add 18 to both sides of the equation 7y - 4 + 18 = 8y - 18 + 18 7y + 14 = 8y Subtract 7y from both sides of the equation 7y - 7y + 14 = 8y - 7y 14 = y ⇒ y = 14

ah42632:

WOW! I couldn´t figure that out if I even tried

jhonyy9:



Montybaby:

C

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CHEERQUEEN We are told that the legs of the isosceles trapezoid are AB and CD AB = 7y - 4 CD = 8y - 18 In any isosceles trapezoid, the two legs are of equal length. Therefore, AB = CD ⇒ 7y - 4 = 8y - 18 Add 18 to both sides of the equation 7y - 4 + 18 = 8y - 18 + 18 7y + 14 = 8y Subtract 7y from both sides of the equation 7y - 7y + 14 = 8y - 7y 14 = y ⇒ y = 14 \(\color{#0cbb34}{\text{End of Quote}}\) good job congrats !